A. Find (a) the mean, (b) the median, (c) the mode, and (d) the midrange.



High Temperatures. The reported high temperatures (in degrees Fahrenheit) for selected world


cities on an October day are shown below. Which measure of central tendency do you think best


describes these data?


62 42


76 38


66 91


79 66


83 77


61 90


62 74


85 63


72 64


64 68


74 42


71

1
Expert's answer
2022-06-06T03:35:43-0400
"62,42,76,38,66, 91,79, 66,83, 77,61, 90,"

"62, 74,85, 63,72, 64,64, 68,74, 42,71"

Arrange in ascending order


"38,42,42,61,62,62,63,64,64,66,66,68,"

"71,72,74,74,76,77,79,83,85,90,91"

(a)


"mean=\dfrac{1}{23}(38+42+42+61+62+62+63"

"+64+64+66+66+68+71+72+74+74"

"+76+77+79+83+85+90+91)"

"=\dfrac{1570}{23}"

(b)


"median=68\ (12^{th} \ place)"

(c)


"mode=62, 42, 64, 66, 74,"

each appeared 2 times


(d)


"midrange=\dfrac{38+91}{2}=64.5"